OWG Meeting 14-Mar-2012

From GlueXWiki
Jump to: navigation, search


  • Announcements
  • Review of minutes from 29-Feb-2012 meeting
  • Software review 7,8-Jun - Mark, David, Elliott
  • TDR - Elliott
  • Cyber Security - Andy Kowalski
  • Hall D Component Inventory Database
  • Counting house - Hovanes
  • Scalers - Dave, Hovanes, Elliott
  • 12GeV planning lines - David, Hovanes, Elliott
  • Elog and Accelerator division projects - Graham
  • DAQ group - Bryan, Dave A, Graham
    • CODA3, EVIO, ECS/CSS, raw event file, disentangler, readout list
  • EPICS - Hovanes, Nerses
  • Controls
  • Beam tests - Alex
  • FDC cosmic ray test - Beni
  • Electronics, trigger and DAQ status reports - Chris, Alex, Dave, Fernando, Ben, Ed, William, Bryan
  • Mantis Task Tracker
  • Midas Elog
  • Calibration/Alignment issues?


1:30 PM Wed 14-Mar-2012 CC F326


Next Meeting

1:30 PM Wed 28-Mar-2012 CC F326


Present: Elliott W, Simon T, Nerses G, Andy Kowalski, Bryan M, Graham H, Vanik K, Hovanes E, Beni Z, David L.


Little is known about this other then we want to have a new version by the Lehman review, late summer or early fall I believe.

Cyber Security

Andy Kowalski from the Computer Center told us about plans for cyber-security in the 12 GeV era.

  • DOE and Mont have imposed strong restrictions concerning access to computers that control equipment in the halls.
  • Additional restrictions will be imposed for sensitive information, probably not relevant for us.
  • Login access to hall computers outside the hall itself from on- or off-site will have to be via two-factor authentication.
  • The preferred two-factor method is to use small portable devices that give you an additional one-time password.
  • Gateways can be used to serve read-only information to computers outside the halls with no special access restrictions.
  • Oscilloscopes and similar devices will live on special networks. They are easily broken into and are often used as "landing points" by attackers.
  • Wireless networks will be expanding to the outdoors in many areas.
  • Linux is not supported on internal laptops due to security concerns. Windows and Mac are ok. Thus Linux laptops must be on the guest network.
  • WPA2 Protected must be used for all wireless connections.
  • Windows tablets are ok on hall networks. IPADs are ok as well.
  • Android devices are totally banned from internal networks, guest only.
  • Users can deploy IPAD apps if desired, but problematic due to Apple policies.
  • New wireless phones will be appearing soon. They connect to wireless Ethernet and thus will work anywhere wireless is available.
  • CorporateTime is being replaced in the next few months with a much better integrated calendar/mail system.

Inventory Database

Elliott briefly demo'd the inventory database. It is a slightly modified clone of the system used by CLAS. Hall D and CLAS plan to share the same code base and further develop the system together. New features coming include power/housing/control hierarchies and the ability to capture cabling information. Elliott will redeploy the system in production mode next week after deleting all data in the test/demo system.

Counting House

Hovanes reported slow progress getting Facilities to run 480V power from the utility room to the rack room for the UPS. Once the UPS is installed we can run power to the rack locations, then install racks and our network switch.

FCC Equipment Specifications

Fernando would prefer that all equipment meet FCC Class B specs, but this is impossible or too expensive in many cases. Fernando would like to be informed when we install any FCC Class A or lower equipment. The big UPS is FCC Class A and it is not clear if we will run UPS power to the hall. The alternative is to purchase FCC Class B UPS's where needed, but they may be expensive and we are told they have to be replaced every two years to remain reliable.


Vanik is working on scaler applications for discriminators for the BCAL test. Alex is working on applications for the FADC scalers.


Graham reported CODA3 fairly close to release, Dave and Carl are reconciling differences in the ROC and EMU concerning how big and little endian encoding is used in the event buffers.


Nerses is continuing work on HV applications for the new CAEN HV mainframes.


Mark S discovered that our Danfysik power supply has a special card that serves out the current readout of the ZFCT, so we don't need to purchase anything (and didn't need the shunt we used in the coil test). Yi is almost ready to purchase the NI PXI system that replaces our original fast/slow coil tap readout system.


Nerses has been looking at recently purchased motion controllers. The goniometer will likely use or be compatible with these controllers.

Beam Tests

Equipment is being tested and moved to Hall B next week.

FDC Test

Beni is chasing down O2 contamination.